There were 476 paper-clips in box A and 52 paper-clips in box B. After an equal number of paper-clips was added into each box, box A had five times as many paper-clips as box B. How many paper-clips were added into box B?
Answer by Paul(988)

Let the number of added clips be x
Box A's clips = 476 --> once added it now becomes: 476+x
Same with box B's clips ---> once added it becomes = 52+x
It says that there were 5 times as many as box B.
Since A has no description; it works with B.
Equation:
5(52+x)=476+x
5x+260=476+x
4x=216
x=54
Hence, about 54 clips were added.
Paul.